What is TIME FOR A LUMP SUM TO DOUBLE

How long will it take $200 to double if it earns the following rates? Compounding occurs once a year. Round each answer to two decimal places.

  1. 6%.

    ? year(s)

  2. 10%.

    ? year(s)

  3. 20%.

    ? year(s)

  4. 100%.

    ? year(s)
Answer #1

In this question, FV = 2 * $200 = $400

a) $400 = $200 * (1 + 6%)n

2 = 1.06n

Taking log of both sides,

LN(2) = n * LN(1.06)

0.693147 = n * 0.058269

n = 11.90 years

b) $400 = $200 * (1 + 10%)n

2 = 1.10n

Taking log of both sides,

LN(2) = n * LN(1.10)

0.693147 = n * 0.09531

n = 7.27 years

c) $400 = $200 * (1 + 20%)n

2 = 1.20n

Taking log of both sides,

LN(2) = n * LN(1.20)

0.693147 = n * 0.182322

n = 3.80 years

d) $400 = $200 * (1 + 100%)n

2 = 1.1n

Taking log of both sides,

LN(2) = n * LN(1.1)

0.693147 = n * 0.693147

n = 1 year